OpenStudy (anonymous):

solve: -5(3x-2)-6x=17

OpenStudy (abb0t):

are you asking me to solve this for you?

OpenStudy (anonymous):

= -1/3??

OpenStudy (anonymous):

did you get - 1/3?

OpenStudy (anonymous):

7/-21 = -1/3??

OpenStudy (abb0t):

no.

OpenStudy (anonymous):

= - 3?

OpenStudy (abb0t):

\(\frac{27}{9}\)

OpenStudy (anonymous):

= - 1/3?

OpenStudy (anonymous):

-9

OpenStudy (abb0t):

27 divided by -9!

OpenStudy (abb0t):

sorry. lol my keyboard sux right now.

OpenStudy (anonymous):

x=-1/3

OpenStudy (anonymous):

-3

OpenStudy (precal):